# Review-Based Cross-Domain Recommendation via Hyperbolic Embedding and Hierarchy-Aware Domain Disentanglement

### v1

@@ 4.2. Feature Extraction
The above word embedding procedure is applied to the source and target domains, respectively. Given word embedding $R^{\mathbb{H}}_{u}$ and $R^{\mathbb{H}}_{i}$ from each domain, we aim to extract useful information. Several feature extraction strategies have been proposed, including domain adaptation (Yuan et al., 2019), variational reconstruction (Liu et al., 2022), personalized transfer (Zhu et al., 2022), contrastive learning (Xie et al., 2022), and domain disentanglement (Peng et al., 2019). Among them, we adopt the domain disentanglement algorithm, which does not obligate user overlapping in both domains (Cao et al., 2022). For this, as illustrated in Figure 2, we employ three types of feature extractors (FEs), all of which consist of simple multi-channel Convolutional Neural Networks (CNNs). Specifically, the shared FE processes datasets from both domains, while the source and target FEs deal with the documents from their domains only (please refer to (Choi et al., 2022) for more details). For the sake of simplicity, we focus on the mechanisms in the source domain since the target domain procedure is the same. The feature extraction process is given by:
@@ 4.3. Hierarchy-Aware Hyperbolic Embedding and Domain Disentanglement
We propose two constraints to achieve domain disentanglement between the extracted features while preserving their hierarchical structure: (1) hierarchy-aware embedding and (2) hierarchy-aware knowledge transfer. The feature extractors (FEs) are configured to preserve such hierarchies during the training as below.
@@
As depicted in Figure 1, one of the remarkable advantages of hyperbolic space is its exponentially expandable capacity. Also, (Khrulkov et al., 2020) reveals that the uncertainty decreases as the embedding gets closer to the boundary of the Poincaré ball. Based on this, HRCF (Yang et al., 2022b) suggests a hyperbolic regularization optimized for the characteristics of a power-law distribution. Specifically, users or items with many interactions (dense) are pulled to the center, while sparse ones are positioned near the boundary. In detail, HRCF identifies the root as the average of the entire embedding and considers it as an origin as follows:
@@
Proof. refer to proof of proposition 4.1 in Section 4.5.
@@
Knowledge transfer has gained substantial attention in the field of cross-domain recommendation. In this regard, we align with the recently proposed domain disentanglement algorithm (Cao et al., 2022; Choi et al., 2022), which operates without the requirement of overlapping users. The fundamental mechanism of domain disentanglement can be delineated as follows: (1) domain-specific features $S,T$ should be readily inferable to their originating domains to mitigate domain discrepancies, and (2) domain-shareable features $\widehat{S},\widehat{T}$ ought to encapsulate domain-indiscriminative information, ensuring pairwise independence between domains (Gretton et al., 2005; Li et al., 2019; Nema et al., 2021). To achieve this, as depicted in Figure 3, we first concatenate the user and item vectors from each feature extractor in the following manner:
@@
Before delving into the scale alignment module which is one of our primary contributions, we introduce the following discussion on previous disentanglement algorithms.
Lemma 4.1 0.
Prior methods of domain disentanglement, such as those proposed by (Peng et al., 2019; Krishnan et al., 2020; Cao et al., 2022; Choi et al., 2022) merely focus on adjusting the scale of the extracted features. In detail, they simply forward these features to the domain discriminator ($F_{d}$) in the following manner:
@@
The $F_{d}$ consists of the two layers of a fully connected neural network and the notation $d_{S},d_{T}\in\{0,1\}$ denotes the predicted domain (0/1 are the source/target). Additionally, $g(\cdot)$ signifies the Gradient Reversal Layer (GRL), which remains inactive during the forward propagation but reverses the sign of the gradient during the back-propagation (for detailed information, refer to (Mansour et al., 2009; Peng et al., 2019)). A major limitation of these methods lies in the disruptive changes in positional information. As elucidated earlier, domain-shareable features tend to converge, while domain-specific features separate apart. Consequently, minor positional changes can cause exponentially magnified effects. A similar issue is observed in other domain alignment methods. To solve the problem, some directly reduce the distance of the same set of users between domains (Zhao et al., 2023), and others leverage variational inference to align the mean and variance of feature distributions (Liu et al., 2022). In this paper, we propose a novel disentanglement strategy that conserves the scale of the extracted information by revising Eq. 18 as follows.
Figure 3. Mechanism of hierarchy-aware domain alignment
@@
We adjust the scale of inputs before applying the domain discriminator, which can preserve the hierarchical structure of features as follows:
@@
Proof. refer to proof of proposition 4.2 in Section 4.5.
Based on this, we can define the loss function of domain discriminator $\mathcal{L}_{d}$ as,
@@
Finally, we claim that scale alignment can also enhance the separability of a discriminator below.
@@
Proof. refer to proof of proposition 4.3 in Section 4.5.
@@ 4.4. Inference and Optimization
Inference. We aim to measure the relativity between a user and an item using their aggregated features. To elaborate, in the center of Figure 2 (source domain), we compute the average (avg) of the outputs from the source and shared FEs as,
(22) $S^{\prime}_{u}={1\over 2}(S_{u}+\widehat{S}_{u}),\,\,\,S^{\prime}_{i}={1\over 2}(S_{i}+\widehat{S}_{i}).$
@@
Finally, we can measure their similarity as below:
@@
where $\mathcal{M}(\cdot)\in[0,1]$ is a MLP with Sigmoid activation function. This adjusts the distance between users and items, $d_{\mathcal{L}}(\cdot)$ (Eq. 4) to reflect the user’s preference for popular items. For optimization, we adopt the hyperbolic margin ($\epsilon=0.1$) ranking loss (Sun et al., 2021) given the positive ($i$) and negative sample ($j$) as below:
@@
Optimization. We define the overall objective function as to minimize the weighted sum of Eq. 16, 21, 26 as,
@@
The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ balance the losses. For each dataset, we find $\lambda_{1}$ and $\lambda_{2}$ through grid search that yielded the best validation score (Fig. 6). The parameter $\theta$ is optimized using the Adam optimizer, with $\delta$ representing the regularization term. Additionally, we practiced early stopping within 300 iterations and applied negative sampling for ratings that meet the condition $y_{u,j}<4$.
@@ 4.5. Theoretical Analysis
Type Method @10 Luxury Beauty All Beauty Digital Music Video Games
@@
Single-Domain NeuMF ND 0.056 0.045 0.062 0.070
HR 0.113 0.089 0.115 0.134
1pt. DeepCoNN ND 0.093 0.089 0.101         0.124
@@
Cross-Domain DAREC ND 0.063 0.048 0.055 0.047 0.043 0.044 0.068 0.073 0.061 0.059 0.066 0.084
HR 0.116 0.101 0.104 0.090 0.082 0.081 0.130 0.140 0.113 0.111 0.127 0.149
1pt. DDTCDR ND 0.072 0.054 0.059 0.054 0.045 0.041 0.065 0.079 0.069 0.062 0.071 0.083
@@
(32) $\mathcal{S}=\phi(Wx)^{T}\phi(Wy)=||x||\cdot||y||\cdot f(x,y)$
4.6. Computational Complexity
In addition to the plain text convolution module ($A$), we employ a hyperbolic Glove that requires a mapping from the hyperbolic space to the Euclidean ones ($B$). Secondly, the discriminator is a simple two-layer neural network and the degree normalization only averages the outputs ($C$). Lastly, the scale alignment has linear complexity as it only matches the magnitudes of the two vectors ($D$). Thus, the complexity is $\mathcal{O}((A+B+C+D)\cdot N_{t})$ which is a linear model proportional to the size of a target domain ($N_{t}$). The implementation of our method is available here [^6] https://anonymous.4open.science/r/HEAD-1F46/README.md.
5. Experiments
We set fundamental questions to provide a comprehensive analysis of the proposed method. Subsequently, we offer details about datasets in Section 5.1, baselines in Section 5.2, and the experimental setup in Section 5.3. The details of the following research questions (RQs) are explained from Section 5.4 to Section 5.7:
5.1. Datasets
Following the prior studies (Choi et al., 2022; Zhao et al., 2023), we evaluate our model and baselines using the benchmark Amazon [^7] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/ 5-core review datasets, which are subsets obtained after excluding the users and items having less than five interactions. We take 12 domain pairs, three as the source domains Clothing (Cloth), CDs and Vinyl (CDs), Toys and Games (Toys), and the other four as the target domains Luxury Beauty, All Beauty, Digital Music, Video Games. We assigned domains to the sources and targets such that the sources have richer interactions than the targets (Ben-David et al., 2010). Also, note that the pairs, (Cloth, Beauty), (CDs, Digital Music), (Toys, Video Games), are relevant to each other. The statistics of the datasets are in Table 2.
5.2. Baselines
For evaluation, we include several state-of-the-art single, cross-domain, and hyperbolic recommendation algorithms. We optimize the single-domain methods only with the target datasets. Moreover, for algorithms (Sun et al., 2021; Wang et al., 2023; Krishnan et al., 2020; Xu and Cai, 2023) that solely rely on rating information, we replaced their embedding mechanism with our review processing module to best ensure fair comparisons.
Single-Domain Approaches: (1) NeuMF (He et al., 2017) is a rating-based probabilistic neural network. (2) DeepCoNN (Zheng et al., 2017) further utilizes reviews as side information in addition to ratings. (3) AHN (Dong et al., 2020) is a review-based hierarchical attention network. (4) HGCF (Sun et al., 2021) suggests margin ranking loss for collaborative filtering in hyperbolic space. (5) HDNR (Wang et al., 2023) alleviates the representation distortion problem through hyperbolic geometry for news recommendation.
Cross-Domain Approaches: (6) DAREC (Yuan et al., 2019) applies domain adaptation to capture domain-shareable rating information. (7) DDTCDR (Li and Tuzhilin, 2020) applies an orthogonal mapping between the same set of users given their rating histories. (8) RC-DFM (Fu et al., 2019) fuses the reviews with rating and transfers knowledge between duplicate users. (9) CATN (Zhao et al., 2020) further employs the features of non-duplicate users with similar interests. (10) MMT (Krishnan et al., 2020) proposes domain-specific and domain-shareable modules without disentanglement. (11) SER (Choi et al., 2022) integrates domain discriminator to disentangle domain-specific and domain-shareable features. (12) DH-GAT (Xu and Cai, 2023) applies graph attention in hyperbolic space for cross-domain entity recognition.
@@ 5.3. Experimental Setup
Following previous studies (Choi et al., 2022; Xu and Cai, 2023; Wang et al., 2023), we split the target domain dataset into 80%/10%/10% for training, validation, and testing without considering temporal sequence. Then, we generated user and item documents excluding interactions present in the validation and test sets. During training, we randomly and iteratively sample from a source domain and from a target domain such that the sizes of two samples are the same. Additionally, we employ an early stopping technique to terminate the training process if the best validation score is not updated for 300 iterations. We show the test accuracy with the best validation score. Finally, the dimension of word embedding is set as 100 for all methods.
@@ 5.4. Model Comparison (RQ1)
Table 4.5 shows the experimental results of the proposed $\mathrm{HEAD}$ and those of the baselines. For quantitative evaluation, we employ Normalized Discounted Cumulative Gain (NDCG@$10$) and Hit Ratio (HR@$10$) as metrics. We analyze the results from two perspectives; (1) the usefulness of reviews and disentangled representation learning, and (2) the comparison between the Euclidean and hyperbolic geometry.
(1) Reviews contribute to mitigating the data sparsity problem, and domain disentanglement also enhances the quality of knowledge transfer by reducing domain discrepancies. Firstly, we observe that methods utilizing only rating information (He et al., 2017; Yuan et al., 2019; Li and Tuzhilin, 2020) show significantly lower performance compared to the models that utilize reviews additionally. This can be due to the small size of the target domain, but we can also presume that user preferences are well reflected in the review information. We also observe that the algorithms perform differently depending on how well the reviews are utilized; attention-based AHN (Dong et al., 2020) significantly outperforms the plain text convolution model, DeepCoNN (Zheng et al., 2017). Additionally, addressing domain discrepancies is also critical to the performance of CDR. For example, MMT (Krishnan et al., 2020), SER (Choi et al., 2022), and our HEAD show stable performance among the CDR techniques regardless of the source domain pairs. This can be inferred that they can separate the domain-shareable and domain-specific knowledge efficiently, making them relatively robust to noise. Additionally, our models exhibit the most stable results, suggesting that scale preservation is helpful in discriminator training, and eventually enables the effective transfer of more useful information to the target domain.
@@
Figure 4. (RQ2) Domain discrimination performance of three types of models; SER (Choi et al., 2022), HEAD (w/o SA), and HEAD (w/ SA). Both figures use the same target domain (Luxury Beauty) with different source domains; (a) Clothing and (b) CDs
Figure 5. (RQ3) We randomly sampled 1,000 items in Digital Music and visualized them based on their degrees. The left figure shows the results without degree normalization in Eq. 14, while the right one represents after applying this method
(2) Hyperbolic embedding achieves better performance compared to Euclidean ones, and HEAD with degree-based normalization and scale adjustment has shown its effectiveness. A notable point is that hyperbolic-based methods HGCF (Sun et al., 2021), and HDNR (Wang et al., 2023) exhibit good performance even without using additional domains. For example, their accuracy is comparable to AHN (Dong et al., 2020) with hierarchical attention mechanism, and DH-GAT (Xu and Cai, 2023) attains the best recommendation quality for some datasets. This is based on the advantages of the vast space in the hyperbolic space, and it is believed to have a positive impact on learning the pairwise distance between the latent representations. In addition to this, our HEAD, which employs degree-based hierarchy correction and scale alignment, achieves a performance improvement of 10.4% compared to SER (Choi et al., 2022). This highlights the benefits of hierarchy alignment in hyperbolic space and the removal of scale information enhances the quality of domain discriminator.

### v3

@@ 4.2. Feature Extraction
The above word embedding procedure is applied to the source and target domains, respectively. Given word embedding $R^{\mathbb{H}}_{u}$ and $R^{\mathbb{H}}_{i}$ from each domain, we aim to extract useful information. Several feature extraction strategies have been proposed, including domain adaptation (Yuan et al., 2019), variational reconstruction (Liu et al., 2022), personalized transfer (Zhu et al., 2022), contrastive learning (Xie et al., 2022), and domain disentanglement (Peng et al., 2019). Among them, we adopt the domain disentanglement algorithm, which does not obligate user overlapping in both domains (Cao et al., 2022a). For this, as illustrated in Figure 2, we employ three types of feature extractors (FEs), all of which consist of simple multi-channel Convolutional Neural Networks (CNNs). Specifically, the shared FE processes datasets from both domains, while the source and target FEs deal with the documents from their domains only (please refer to (Choi et al., 2022) for more details). For the sake of simplicity, we focus on the mechanisms in the source domain since the target domain procedure is the same. The feature extraction process is given by:
@@ 4.3. Hierarchy-Aware Hyperbolic Embedding and Domain Disentanglement
We propose two constraints to achieve domain disentanglement between the extracted features while preserving the hierarchy.
@@
Recent work (Khrulkov et al., 2020) reveals that the uncertainty decreases as the embedding gets closer to the boundary of the Poincaré ball. HRCF (Yang et al., 2022b) suggests a hyperbolic regularization optimized for the characteristics of a power-law distribution. Specifically, users or items with many interactions (dense) are pulled to the center, while sparse ones are placed near the boundary. For this, HRCF identifies the root as the average of the entire embedding to make it as an origin below:
@@
Proof. see proof of proposition 4.1 in Section 4.6.
@@
Knowledge transfer has gained substantial attention in the field of cross-domain recommendation. In this regard, we align with the recently proposed domain disentanglement algorithm (Cao et al., 2022a; Choi et al., 2022), which operates without the requirement of overlapping users. The fundamental mechanism of domain disentanglement can be delineated as follows: (1) domain-specific features $S,T$ should be readily inferable to their originating domains to mitigate domain discrepancies, and (2) domain-shareable features $\widehat{S},\widehat{T}$ ought to encapsulate domain-indiscriminative information, ensuring pairwise independence between domains (Gretton et al., 2005; Li et al., 2019; Nema et al., 2021). For this, we first concatenate the user and item vectors from each feature extractor as below:
@@
Before delving into the scale alignment module, we introduce the following discussion on previous disentanglement algorithms. Prior methods of domain disentanglement, such as those proposed by (Peng et al., 2019; Krishnan et al., 2020; Cao et al., 2022a; Choi et al., 2022; Zhang et al., 2022) merely focus on preserving the scale of the extracted features. In detail, they simply forward these features to the domain discriminator ($F_{d}$) in the following manner:
@@
The $F_{d}$ consists of the two layers of a fully connected neural network and the notation $d_{S},d_{T}\in\{0,1\}$ denotes the predicted domain (0/1 are the source/target). Additionally, $g(\cdot)$ signifies the Gradient Reversal Layer (GRL), which remains inactive during the forward propagation but reverses the sign of the gradient during the back-propagation (for detailed information, refer to (Mansour et al., 2009; Peng et al., 2019)). A major limitation of these methods lies in the disruptive changes in positional information. As elucidated earlier, domain-shareable features tend to converge, while domain-specific features separate apart. Consequently, minor positional changes can cause exponentially magnified effects. A similar issue is observed in other domain alignment methods. To solve the problem, some directly reduce the distance of the same set of users between domains (Zhao et al., 2023a), and others leverage variational inference to align the mean and variance of feature distributions (Liu et al., 2022). In this paper, we propose a novel disentanglement strategy that conserves the scale of the extracted information by revising Eq. 18 as follows.
@@
We adjust the scale of inputs before applying the domain discriminator as follows:
@@
Based on this, we can define the domain loss $\mathcal{L}_{d}$ as,
@@
Proof. see proof of proposition 4.2 in Section 4.6.
Finally, we claim that scale alignment can also enhance the separability of a discriminator in the proposition below.
@@
Proof. see proof of proposition 4.3 in Section 4.6.
@@ 4.4. Inference and Optimization
Inference. We aim to measure the relativity between a user and an item using their aggregated features. To elaborate, in the middle of Figure 2 (source domain), we compute the average (avg) of the outputs from the source ($S$) and shared FEs ($\widehat{S}$), adding the latent of user-item interaction vectors ($p_{u}$ and $p_{i}$) as follows:
(22) $S^{\prime}_{u}={1\over 2}(S_{u}+\widehat{S}_{u})+p_{u},\,\,\,S^{\prime}_{i}={1\over 2}(S_{i}+\widehat{S}_{i})+p_{i}$
@@
Finally, we can measure their distance as below:
@@
where $\mathcal{M}(\cdot)\in[0,1]$ is a MLP with Sigmoid activation function. This adjusts the distance between users and items, $d_{\mathcal{L}}(\cdot)$ (Eq. 4) to reflect the user’s preference for popular items. For optimization, we adopt the hyperbolic margin ($\epsilon=0.1$) ranking loss (Sun et al., 2021) given the positive ($i$) and negative sample ($j$) as,
@@
Optimization. We define the overall objective function as to minimize the weighted sum of Eq. 16, 21, 26 as below:
@@
The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ balance the losses. For each dataset, we find $\lambda_{1}$ and $\lambda_{2}$ through grid search that yielded the best validation score (Fig. 5). The parameter $\theta$ is optimized using the Adam optimizer, with $\delta$ representing the regularization term. Additionally, we practiced early stopping within 300 iterations and applied negative sampling for ratings that meet the condition of $y_{u,j}\leq 3$.
4.5. Time Complexity
In addition to the plain text convolution module ($A$), we employ a hyperbolic Glove that requires a mapping from the hyperbolic space to the Euclidean ones ($B$). Secondly, the discriminator is a simple two-layer neural network and the degree normalization only averages the outputs ($C$). Lastly, the scale alignment has linear complexity as it only matches the magnitudes of the two vectors ($D$). Thus, the complexity is $\mathcal{O}((A+B+C+D)\cdot N_{t})\approx\mathcal{O}(N_{t})$, which is a linear model proportional to the size of a target domain.
@@ 4.6. Theoretical Analysis
(32) $\mathcal{S}=\phi(Wx)^{T}\phi(Wy)=||x||\cdot||y||\cdot f(x,y)$
5. Experiments
We set fundamental questions to provide a comprehensive analysis of the proposed method. The details of the following research questions (RQs) are explained from Section 5.2 to Section 5.5:
@@ 5. Experiments
Method @10 Luxury Beauty All Beauty Digital Music Video Games
@@
Single-Domain DeepCoNN ND 0.093 0.089 0.101 0.124
@@
Cross-Domain DDTCDR ND 0.072 0.054 0.059 0.054 0.045 0.041 0.065 0.079 0.069 0.062 0.071 0.083
@@ 5.1. Experimental Setup
Following the prior studies (Choi et al., 2022; Zhao et al., 2023a), we evaluate our model using Amazon [^7] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/ 5-core review datasets. As shown in Table 1, we take 12 domain pairs, three as the source with richer interactions and four as the target with sparse ones (Ben-David et al., 2010). The pairs of (Cloth, Beauty), (CDs, Digital Music), (Toys, Video Games) are relevant to each other. The target domain datasets are split into 80%/10%/10% for training, validation, and testing without considering a temporal sequence same as (Choi et al., 2022; Xu and Cai, 2023; Wang et al., 2023). Additionally, we employ an early stopping technique to terminate the training process if the best validation score is not updated for 300 iterations. The dimension of word embedding is set as 100 for all methods.
@@ 5.2. Model Comparison (RQ1)
In Table 5, we show the results using Normalized Discounted Cumulative Gain (NDCG@$10$) and Hit Ratio (HR@$10$).
(1) Reviews and domain disentanglement enhance the quality of recommendation. Firstly, we observe that methods utilizing only rating information (He et al., 2017; Yuan et al., 2019; Li and Tuzhilin, 2020) show significantly lower performance compared to the review-based models. This can be due to the small size of the target domain, but we can also presume that user preferences are well reflected in the review information. We also observe that the algorithms perform differently depending on how well the reviews are utilized; attention-based AHN (Dong et al., 2020) significantly outperforms the plain text convolution model, DeepCoNN (Zheng et al., 2017). Additionally, addressing domain discrepancies is also critical to the performance of CDR. For example, MMT (Krishnan et al., 2020), SER (Choi et al., 2022), and our HEAD show stable performance among the CDR techniques regardless of the source domain pairs. This can be inferred that they can separate the domain-shareable and domain-specific knowledge efficiently, making them relatively robust to noise. Additionally, our models exhibit the most stable results, suggesting that scale preservation is helpful in discriminator training.
(2) Hyperbolic embedding achieves better performance compared to Euclidean ones, and HEAD with degree-based normalization and scale adjustment has shown its effectiveness. A notable point is that hyperbolic-based methods HGCF (Sun et al., 2021), GDCF (Zhang et al., 2022), and HDNR (Wang et al., 2023) exhibit good performance even without using additional domains. For example, their accuracy is comparable to AHN (Dong et al., 2020) with hierarchical attention mechanism, and DH-GATH (Xu and Cai, 2023) attains the best recommendation quality for some datasets. This is based on the advantages of the vast space in the hyperbolic space, which has a positive impact on learning the pairwise distance between the latent representations. In addition to this, our HEAD, which employs degree-based hierarchy correction and scale alignment, achieves a performance improvement of 10.4% compared to SER (Choi et al., 2022). This highlights the benefits of hierarchy alignment in hyperbolic space and the removal of scale information enhances the quality of the domain discriminator.
@@
Figure 3. (RQ2) Domain discrimination performance of three methods with similar and dissimilar domain pairs
Figure 4. (RQ3) We randomly sampled 1,000 items in Digital Music and visualized them based on their degrees